Mickey is eight years younger than Minnie. In one year, Minnie will be three times as old as Mickey was five years ago. Find their ages now.
Answer by ankor@dixie-net.com(22740)

Let m = Mickey's age
let n = Minnie's age
:
Write an equation for each statement, simplify as much as possible.
:
Mickey is eight years younger than Minnie.
m = n - 8
or
n = m + 8
:
In one year, Minnie will be three times as old as Mickey was five years ago.
n + 1 = 3(m-5)
n + 1 = 3m - 15
n = 3m - 15 -1
n = 3m - 16
replace n with (m+8),from the 1st statement
m + 8 = 3m - 16
8 + 16 = 3m - m
24 = 2m
m = 24/2
m = 12 yrs is Mickey's age
:
I'll let you find Minnie's age, check your solutions in the 2nd statement